# VITA: Towards Open-Source Interactive Omni Multimodal LLM

### v1

@@ Abstract
The remarkable multimodal capabilities and interactive experience of GPT-4o underscore their necessity in practical applications, yet open-source models rarely excel in both areas. In this paper, we introduce VITA, the first-ever open-source Multimodal Large Language Model (MLLM) adept at simultaneous processing and analysis of Video, Image, Text, and Audio modalities, and meanwhile has an advanced multimodal interactive experience. Starting from Mixtral 87B as a language foundation, we expand its Chinese vocabulary followed by bilingual instruction tuning. We further endow the language model with visual and audio capabilities through two-stage multi-task learning of multimodal alignment and instruction tuning. VITA demonstrates robust foundational capabilities of multilingual, vision, and audio understanding, as evidenced by its strong performance across a range of both unimodal and multimodal benchmarks. Beyond foundational capabilities, we have made considerable progress in enhancing the natural multimodal human-computer interaction experience. To the best of our knowledge, we are the first to exploit non-awakening interaction and audio interrupt in MLLM. We design additional state tokens, and corresponding training data and strategies to perceive various interaction scenarios. The deployment of VITA employs a duplex scheme, where one model is responsible for generating responses to user queries, and the other continuously tracks environmental inputs, selectively outputting new responses with updated interactions. This allows VITA to feature impressive human-machine interaction functionalities such as non-awakening interaction and audio interrupt interaction. VITA is the first step for the open-source community to explore the seamless integration of multimodal understanding and interaction. While there is still lots of work to be done on VITA to get close to close-source counterparts, we hope that its role as a pioneer can serve as a cornerstone for subsequent research.
@@ 3.4.2 Audio Interrupt Interaction
To achieve this, we propose the duplex deployment framework. As illustrated in Fig. 1, two VITA models are deployed concurrently. Under a typical condition, the Generation model answers user queries. Simultaneously, the Monitoring model detects environmental sounds during the generation process. It disregards non-query user sounds, i.e., noisy audio, but ceases the Generation model’s progress when it identifies query audio. The Monitoring model subsequently consolidates the historical context and responds to the latest user query. At this point, the identities of the Generation model and the Monitoring model are transformed.
@@ 4 Evaluation
Test_Net | Test_Meeting | Dev_clean | Dev_other | est_clean | est_other | |||
@@
Multimodal Performance. To assess multimodal capabilities, we evaluate VITA on four representative benchmarks, including MME [13], OCRBench [32], HallusionBench [16], and Video-MME [14]. As depicted in Fig. 5, in terms of image understanding, VITA outperforms image specialized open-source model LLaVA-Next [30] and is close to closed-source model Gemini 1.5 Pro [44]. In video understanding, VITA surpasses video specialized open-source model Video-CCAM999https://github.com/QQ-MM/Video-CCAM. Although there is a gap between VITA and the video-specialized LLaVA-Next-Video [57], this is acceptable given that VITA supports a broader range of modalities and prioritizes interaction. However, it is worth noting that a substantial gap still exists between current open-source models and proprietary models in terms of video understanding capabilities.

### v2

@@ Abstract
The remarkable multimodal capabilities and interactive experience of GPT-4o underscore their necessity in practical applications, yet open-source models rarely excel in both areas. In this paper, we introduce VITA, the first-ever open-source Multimodal Large Language Model (MLLM) adept at simultaneous processing and analysis of Video, Image, Text, and Audio modalities, and meanwhile has an advanced multimodal interactive experience. Starting from Mixtral 87B as a language foundation, we expand its Chinese vocabulary followed by bilingual instruction tuning. We further endow the language model with visual and audio capabilities through two-stage multi-task learning of multimodal alignment and instruction tuning. VITA demonstrates robust foundational capabilities of multilingual, vision, and audio understanding, as evidenced by its strong performance across a range of both unimodal and multimodal benchmarks. Beyond foundational capabilities, we have made considerable progress in enhancing the natural multimodal human-computer interaction experience. We design additional state tokens, and corresponding training data and strategies to perceive various interaction scenarios. The deployment of VITA employs a duplex scheme, where one model is responsible for generating responses to user queries, and the other continuously tracks environmental inputs, selectively outputting new responses with updated interactions. This allows VITA to feature impressive human-machine interaction functionalities such as non-awakening interaction and audio interrupt interaction. VITA is the first step for the open-source community to explore the seamless integration of multimodal understanding and interaction. While there is still lots of work to be done on VITA to get close to close-source counterparts, we hope that its role as a pioneer can serve as a cornerstone for subsequent research.
@@ 3.4.2 Audio Interrupt Interaction
To achieve this, we propose the duplex deployment framework, which is also an important research direction in audio field [39]. As illustrated in Fig. 1, two VITA models are deployed concurrently. Under a typical condition, the Generation model answers user queries. Simultaneously, the Monitoring model detects environmental sounds during the generation process. It disregards non-query user sounds, i.e., noisy audio, but ceases the Generation model’s progress when it identifies query audio. The Monitoring model subsequently consolidates the historical context and responds to the latest user query. At this point, the identities of the Generation model and the Monitoring model are transformed.
@@ 4 Evaluation
Test_Net | Test_Meeting | Dev_clean | Dev_other | Test_clean | Test_other | |||
@@
Multimodal Performance. To assess multimodal capabilities, we evaluate VITA on ten representative benchmarks, including MME [14], MMBench [34], MMStar [6], MMMU [59], MathVista [37], HallusionBench [17], AI2D [20], OCRBench [35], MMVet [58], and Video-MME [15]. The last one is a video benchmark, and the others belong to image evaluation sets. As depicted in Fig. 5, in terms of image understanding, VITA shows comparable performance with image specialized open-source model LLaVA-Next [32] and is close to closed-source model Gemini 1.5 Pro [49]. In video understanding, although there is a small gap between VITA and the video-specialized LLaVA-Next-Video [64], this is acceptable given that VITA supports a broader range of modalities and prioritizes interaction. However, it is worth noting that a substantial gap still exists between current open-source models and proprietary models.